# Suppression of Bogoliubov momentum pairing and emergence of non-Gaussian correlations in ultracold interacting Bose gases

### v1

@@ Acknowledgements
We thank E. Gradova for her contribution to the bootstrap analysis, T. Chalopin for a careful reading of the manuscript, D. Boiron and the members of the Quantum Gas group at Institut d’Optique for insightful discussions. We acknowledge financial support from the Région Ile-de-France in the framework of the DIM SIRTEQ, the “Fondation d’entreprise iXcore pour la Recherche”, the French National Research Agency (Grant number ANR-17-CE30-0020-01) and France 2030 programs of the French National Research Agency (Grant number ANR-22-PETQ-0004). All numerical simulations were made on the PSMN cluster at the ENS of Lyon.
@@ Methods
Measurement of connected correlations at opposite momenta. As indicated in the main text, the volume $\Omega_{k}$ of the depletion in which we compute correlations is the volume between two cubes of sides $2\times 0.3~{}k_{d}$ and $2\times 0.7~{}k_{d}$. Since $\Omega_{k}$ corresponds to $1.8\leq|k|\xi\leq 3.0$ at $U/J=2$ and $1.2\leq|k|\xi\leq 2.0$ at $U/J=22$, with $\xi$ the healing length, it excludes by design the phononic part $\left(|k|\xi\ll 1\right)$ of the thermal depletion. Two-body correlations at opposite momenta are computed from (i) calculating the histogram of momentum sums ${\bm{k}}_{1}+{\bm{k}}_{2}$ for all pairs of atoms detected in a run of the experiment (from which we subtract the product of average occupations $\langle N(\bm{k}_{1})\rangle\langle N(\bm{k}_{2})\rangle$), and (ii) integrating this histogram over a cubic volume centered on ${\bm{k}}_{1}+{\bm{k}}_{2}={\bm{0}}$ and of size $0.18~{}k_{d}$ set by the width of the two-body correlation function.
Comparison of numerical results from quantum rotor model with experiments. Connected correlations $\overline{G}_{\rm c}^{(2)}(0)$ and $\overline{G}_{\rm c}^{(3)}(0)$ are computed numerically with the quantum rotor model over a cubic corona $\Omega_{k}$ between two cubes of sides $(2\times 0.3~{}k_{d}$ and $2\times 0.5~{}k_{d}$. These calculations are performed with a large, integer filling $n_{\rm QR}$ of the 3D lattice (Supplemental material), which leaves us with an adjustable parameter to compare with experiments (performed in the Bose-Hubbard regime with one atom per site at the trap center). The numerically-computed population $\langle N_{\Omega_{k}}^{th}\rangle$ of the modes in momentum space changes with the filling $n_{\rm QR}$. To plot the numerical results of $\overline{G}_{\rm c}^{(2)}(0)$ (resp. $\overline{G}_{\rm c}^{(3)}(0)$) in the main text, we have rescaled their amplitude by $(\langle N_{\Omega_{k}}\rangle/\langle N_{\Omega_{k}}^{th}\rangle)^{2}$ (resp. $(\langle N_{\Omega_{k}}\rangle/\langle N_{\Omega_{k}}^{th}\rangle)^{3}$), where $\langle N_{\Omega_{k}}\rangle$ is the measured momentum population observed in the experiment.
@@
I Normalized two-body correlation functions
We evaluate the normalized two-body correlation function $g^{(2)}(\bm{k},\bm{k}^{\prime})$ in momentum space in the depletion of interacting Bose gases (in the volume $\Omega_{k}$ shown in Fig. 2a of the main text).
In the weakly-interacting regime, the two-body correlation function $g^{(2)}(\bm{k},\bm{k}^{\prime})$ exhibits two peaks Tenart et al. (2021): one located at $\bm{k}^{\prime}=\bm{k}$ due to bosonic bunching (also coined Hanbury-Brown and Twiss (HBT) correlations) and one located at $\bm{k}^{\prime}=-\bm{k}$ due to interaction-induced quantum fluctuations. We fit these peaks with Gaussian functions to obtain their amplitudes, $g_{\rm HBT}^{(2)}(0)$ and $g_{\rm pairs}^{(2)}(0)$, with the results plotted in Fig. S1.
Figure S1: Magnitudes $g_{\rm HBT}^{(2)}(0)$ of local $k/k$ two-body correlations (bunching) and $g_{\rm pairs}^{(2)}(0)$ of normalized two-body correlations at opposite momenta $k/-k$ plotted as a function of the interaction strength $u$ for data sets shown in Fig. 2 of the main text. Horizontal and vertical error bars correspond to one standard deviation.
The bunching amplitude is perfectly-contrasted in all regimes of interaction, $g_{\rm HBT}^{(2)}(0)=2.00(5)$, extending previous observations Cayla et al. (2020) to the strongly-interacting regime. In contrast, $g_{\rm pairs}^{(2)}(0)$ is found to decrease with the interaction $u$. This decrease is compatible with the one expected in Bogoliubov theory: the magnitude $g_{\rm pairs}^{(2)}(0)-1$ is indeed expected to decrease with the inverse of the mode population $1/N(\bm{k})$, a property of two-mode squeezed states Tenart et al. (2021). No qualitative change is expected when the gas enters the strongly correlated regime where the Bogoliubov approximation fails: $g_{\rm pairs}^{(2)}(0)$ also decreases when the population of two-mode squeezed states decreases. This illustrates that normalized correlation functions are not suited to study the hierarchy of momentum-correlated subsets introduced in the main text (see Fig. 1a) and it motivates the use of connected correlations to reveal the correlated subsets of $n$ modes.
II Extracting two-body connected correlations
In finite-size systems, correlation functions have a finite extent – the correlation volume – whose volume is related to that of one mode. In our experiment, the normalized correlation function $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$ is expected to be isotropic due to the presence of an isotropic harmonic trap and is found, empirically, to be well fitted by an isotropic Gaussian function. We introduce the correlation length $l_{c}$ as the size of $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$ at $1/e^{2}$:
$g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})=1+\frac{A\sqrt{2}}{\sqrt{\pi}l_{c}}e^{-\frac{2\delta\bm{k}^{2}}{l_{c}^{2}}}.$
Thus 95% of atom pairs are expected to be found in the spherical volume where $|\delta\bm{k}|\in[0,l_{c}]$.
@@ II Extracting two-body connected correlations
The two-body connected correlation $G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ is obtained from integrating $G^{(2)}_{c}({\bm{k}},-{\bm{k}}+\delta\bm{k})$ over a cubic volume of size $\Delta k$:
$\begin{split}G^{(2)}_{c}({\bm{k}},-{\bm{k}})=\int_{[-\Delta k/2,\Delta k/2]^{3}}d\delta\bm{k}\ G^{(2)}_{c}(\bm{k},-\bm{k}+\delta\bm{k})).\end{split}$ (S1)
Since $G^{(2)}_{c}({\bm{k}},-{\bm{k}}+\delta\bm{k})\propto g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})-1$ it follows that $G^{(2)}_{c}({\bm{k}},-{\bm{k}})\propto\left[\mathrm{erf}\left(\frac{\Delta k}{\sqrt{2}l_{c}}\right)\right]^{3}$. We choose the volume $\Delta k^{3}$ such that the above integral on $d\delta{\bm{k}}$ saturates: we set $\Delta k=0.18\ k_{d}\simeq 3\times l_{c}$ such that $G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ accounts for 99% of the pairing amplitude, theoretically. At moderate interaction strengths $U/J<20$, this is verified in the experiment, as illustrated in Fig. S2 where we plot the integrated version $\bar{G}^{(2)}_{c}(\bm{0})=\sum_{\bm{k}}G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ normalized by the integration volume $\Delta k^{3}$. We find indeed that $\bar{G}^{(2)}_{c}(\bm{0})/\Delta k^{3}$ decays to zero over the expected range $\Delta k\sim 0.18\ k_{d}$, with the variation $\left[\mathrm{erf}\left(\frac{\Delta k}{\sqrt{2}l_{c}}\right)\right]^{3}/\Delta k^{3}$ obtained from the bell-shaped function of the normalized correlations. This is reminiscent of the fact that we observe a peak in the normalized correlation $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$ centered on $\delta\bm{k}={\bm{0}}$ for $U/J<20$.
@@
In contrast, at $U/J=20$ (and similarly at $U/J=22$), $\bar{G}^{(2)}_{c}(\bm{0})/\Delta k^{3}$ follows a very different trend. While it takes slightly positive values at vanishingly small values of $\Delta k$, it is constant and negative otherwise. This different behaviour is at the origin of the negative values $\bar{G}^{(2)}_{c}(\bm{0})$ plotted in Fig. 2 of the main text. Observing $\bar{G}^{(2)}_{c}(\bm{0})<0$ implies that the number of coincidence counts in modes at opposite momenta is smaller than the coincidence counts expected in an uncorrelated ensemble of atoms with a similar momentum density $N({\bm{k}})=\langle a^{\dagger}_{\bm{k}}a_{\bm{k}}\rangle$, and hence that modes at opposite momenta are anti-correlated.
III Connected correlations $G^{(2)}_{c}({\bf 0})$ and $\langle a^{\dagger}_{\bf k}a^{\dagger}_{-\bf k}a_{\bf k}a_{-{\bf k}}\rangle_{c}$
We discuss the relation between the connected two-body correlations $G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ and connected four-operator correlations $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}$ in the condensate depletion.
In the condensate depletion and in a number-conserving approach, the connected correlations of four operators write,
@@
-- $|\langle a^{\dagger}_{\bm{k}}a_{-\bm{k}}\rangle|^{2}-\langle a^{\dagger}_{\bm{k}}a_{\bm{k}}\rangle\langle a^{\dagger}_{-\bm{k}}a_{-\bm{k}}\rangle,$
and, in turn, $G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ can be expressed as,

### v2

@@ Acknowledgements
We thank E. Gradova for her contribution to the bootstrap analysis, T. Chalopin for a careful reading of the manuscript, D. Boiron and the members of the Quantum Gas group at Institut d’Optique for insightful discussions. We acknowledge financial support from the Région Ile-de-France in the framework of the DIM SIRTEQ, the “Fondation d’entreprise iXcore pour la Recherche”, the French National Research Agency (Grant number ANR-17-CE30-0020-01) and France 2030 programs of the French National Research Agency (Grant numbers ANR-22-PETQ-0004 and ANR-11-IDEX-0003). All numerical simulations were made on the PSMN cluster at the ENS of Lyon.
@@ Methods
Measurement of connected correlations at opposite momenta. The two-body correlations shown in the main text are calculated in the depletion of the condensates, considering atoms belonging to the volume $\Omega_{k}$ defined as the volume between two cubes of sides $2\times 0.3~{}k_{d}$ and $2\times 0.7~{}k_{d}$. Since $\Omega_{k}$ corresponds to $1.8\leq|k|\xi\leq 3.0$ at $U/J=2$ and $1.2\leq|k|\xi\leq 2.0$ at $U/J=22$, with $\xi$ the healing length, it excludes by design the phononic part $\left(|k|\xi\ll 1\right)$ of the thermal depletion. Two-body correlations at opposite momenta $G_{c}^{(2)}$ are computed from (i) calculating the histogram of momentum sums ${\bm{k}}_{1}+{\bm{k}}_{2}$ for all pairs of atoms detected in $\Omega_{k}$ in a run of the experiment Tenart et al. (2021) (from which we subtract the product of average occupations $\langle N(\bm{k}_{1})\rangle\langle N(\bm{k}_{2})\rangle$), and (ii) integrating this histogram over a cubic volume centered on ${\bm{k}}_{1}+{\bm{k}}_{2}={\bm{0}}$ and of size $0.18~{}k_{d}$ set by the width of the two-body correlation function.
Comparison of numerical results from quantum rotor model with experiments. Connected correlations $\overline{G}_{\rm c}^{(2)}(0)$ and $\overline{G}_{\rm c}^{(3)}(0)$ are computed numerically with the quantum rotor model over a cubic corona $\Omega_{k}$ between two cubes of sides $2\times 0.3~{}k_{d}$ and $2\times 0.5~{}k_{d}$. These calculations are performed with a large, integer filling $n_{\rm QR}$ of the 3D lattice (Supplemental material), which leaves us with an adjustable parameter to compare with experiments (performed in the Bose-Hubbard regime with one atom per site at the trap center). The numerically-computed population $\langle N_{\Omega_{k}}^{th}\rangle$ of the modes in momentum space changes with the filling $n_{\rm QR}$. To plot the numerical results of $\overline{G}_{\rm c}^{(2)}(0)$ (resp. $\overline{G}_{\rm c}^{(3)}(0)$) in the main text, we have rescaled their amplitude by $(\langle N_{\Omega_{k}}\rangle/\langle N_{\Omega_{k}}^{th}\rangle)^{2}$ (resp. $(\langle N_{\Omega_{k}}\rangle/\langle N_{\Omega_{k}}^{th}\rangle)^{3}$), where $\langle N_{\Omega_{k}}\rangle$ is the measured momentum population observed in the experiment.
@@
I Thermometry of the lattice Bose gases
To estimate the temperature of the lattice Bose gases realised in the experiment, we compare the measured momentum densities $\rho({\bm{k}})$ to ab-initio calculations for the Bose-Hubbard hamiltonian (BHH). More specifically, we perform unbiased Quantum Monte-Carlo calculations of the momentum density $\rho_{\rm QMC}({\bm{k}},T)$ for the BHH model using all the calibrated experimental parameters (particle number, trapping potentials), except for the temperature $T$ which is not measured in the experiment. At each value of $U/J$, we minimise the distance between $\rho({\bm{k}})$ and $\rho_{\rm QMC}({\bm{k}},T)$ to estimate the temperature $T$ in the experiment. This procedure is described in Carcy et al. (2021).
For the data shown in this work, the estimated reduced temperature varies smoothly from $T/J=0.8(3)$ at $U/J=1$ to $T/J=3.1(2)$ at $U/J=22$. This increase is compatible with an isentropic variation of $U/J$ Carcy et al. (2021), as stated in the main text.
II Two-body connected correlations in the experiment
From the measured 3D distributions of individual atoms in momentum space, we calculate the connected two-body correlations $G_{c}^{(2)}({\bm{k}},{\bm{k}}^{\prime})=\langle N({\bm{k}})N({\bm{k}}^{\prime})\rangle-\langle N({\bm{k}})\rangle\langle N({\bm{k}}^{\prime})\rangle$. More specifically, we evaluate the terms $\langle N({\bm{k}})N({\bm{k}}^{\prime})\rangle$ and $\langle N({\bm{k}})\rangle\langle N({\bm{k}}^{\prime})\rangle$ independently in the depletion of interacting Bose gases (in the volume $\Omega_{k}$ shown in Fig. 2a of the main text) proceeding similarly to the method introduced in Tenart et al. (2021). In short, we first consider all the couples of detected atoms $({\bm{k}},{\bm{k}}^{\prime})$ in a given shot and fill a 3D histogram with the sum of their momenta, ${\bm{K}}={\bm{k}}+{\bm{k}}^{\prime}$. Here the elementary volume of the histogram is a cube of size $0.012k_{d}$. The term $\langle N({\bm{k}})N({\bm{k}}^{\prime})\rangle$ is obtained from averaging the 3D histogram over all the shots. The second term, $\langle N({\bm{k}})\rangle\langle N({\bm{k}}^{\prime})\rangle$, is obtained from first merging the atoms from all shots together and subsequently computing the 3D histograms of their sum ${\bm{K}}$.
Atoms with opposite momenta populate the ${\bm{K}}\simeq{\bm{0}}$ region of the histogram. Considering an elementary volume located at a distance $\delta{\bm{k}}$ from ${\bm{K}}={\bm{0}}$ provides us with the value of $G^{(2)}_{c}({\bm{K}}=\delta\bm{k})=\sum_{\bm{k}}G^{(2)}_{c}({\bm{k}},-{\bm{k}}+\delta\bm{k})$. In addition, we eliminate an artifact due to some ambiguity in the reconstruction of the 3D momenta of the atoms from the timestamps of the raw detection events. This artifact occurs at low values of $U/J$ for which the density of the condensate peak located at ${\bm{k}}={\bm{0}}$ is high enough to induce this effect that would indeed contribute to the connected correlations at opposite momenta (see Ténart (2021) for further details).
@@ II Two-body connected correlations in the experiment
The two-body connected correlation $\bar{G}^{(2)}_{c}(\mathbf{0})$ is then obtained from integrating $G^{(2)}_{c}(\bm{K}=\delta\bm{k})$ over a cubic volume of size $\Delta k$:
@@
In the experiment, the normalised correlation function $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$ is expected to be isotropic due to the presence of an isotropic harmonic trap (of frequency $\omega/2\pi=140(5)\sqrt{s}~{}$z where $s$ the lattice amplitude in recoil energy). In the regime of moderate interactions where a peak is indeed observed in $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$, this peak is empirically found to be well fitted by an isotropic Gaussian function. We introduce the correlation length $l_{c}$ as the size of $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$ at $1/e^{2}$:
$g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})=1+\frac{A\sqrt{2}}{\sqrt{\pi}l_{c}}e^{-\frac{2\delta\bm{k}^{2}}{l_{c}^{2}}}.$ (S1)
Thus 95% of atom pairs are expected to be found in the spherical volume where $|\delta\bm{k}|\in[0,l_{c}]$. Since $G^{(2)}_{c}(\bm{K}=\delta\bm{k})\propto\sum_{\bm{k}}\ g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})-1$ it follows from Eq. (S1) that
$\bar{G}^{(2)}_{c}(\bm{0})\propto\left[\mathrm{erf}\left(\frac{\Delta k}{\sqrt{2}l_{c}}\right)\right]^{3}.$ (S2)
Figure S1: Two-body connected correlation $\bar{G}^{(2)}_{c}(\bm{0})$ as a function of the size of the integration volume $\Delta k$, from $U/J=1$ through $12.5$. The error bars correspond to 68%-confidence intervals obtained from a bootstrap method. The grey solid line is the empirical model following Eq. (S2) associated with the correlation length $l_{c}$ of the normalised two-body correlation $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$. The orange data correspond to an integration of $\Delta k=0.18\ k_{d}$ and are the ones shown in the main text. The black dashed lines indicate the integration corresponding to three times the correlation length $l_{c}$.
At moderate interaction strengths $U/J<20$, this is verified in the experiment, as illustrated in Fig. S1 where we plot $\bar{G}^{(2)}_{c}(\bm{0})$ as a function of $\Delta k$. We find indeed that $\bar{G}^{(2)}_{c}(\bm{0})$ increases and saturates at the expected range $\Delta k\simeq 3\times l_{c}\sim 0.18\ k_{d}$, with the variation $\left[\mathrm{erf}\left(\frac{\Delta k}{\sqrt{2}l_{c}}\right)\right]^{3}$ obtained from eq. (S2). This is reminiscent of the fact that we observe a peak in the normalised correlation $g^{(2)}(\bm{k},-\bm{k}+\delta\bm{k})$ centered on $\delta\bm{k}={\bm{0}}$ for $U/J<20$. To plot the amplitudes of connected correlation in the main text, we choose the volume $\Delta k^{3}$ such that the above integral on $d\delta{\bm{k}}$ saturates: we set $\Delta k=0.18\ k_{d}$ such that $\bar{G}^{(2)}_{c}(\bm{0})$ accounts for 99% of the pairing amplitude, theoretically. At higher $\Delta k$ only uncorrelated atoms are being included in $\bar{G}^{(2)}_{c}(\bm{0})$ and thus its value barely changes within error bars. However, since uncorrelated atoms contribute (equally) to the two terms in $G_{c}^{(2)}({\bm{k}},{\bm{k}}^{\prime})$, the error bars increase as a consequence of $G_{c}^{(2)}({\bm{k}},{\bm{k}}^{\prime})$ being a (constant) difference of increasingly large numbers.
A complementary visualisation consists in plotting an intensive quantity: $\bar{G}^{(2)}_{c}(\bm{0})$ normalised by the integration volume $\Delta k^{3}$ as shown in Fig. S2. We find that $\bar{G}^{(2)}_{c}(\bm{0})/\Delta k^{3}$ decays to zero over the expected range $\Delta k\simeq 3\times l_{c}\sim 0.18\ k_{d}$ with the expected variation.
@@
In contrast, at $U/J=20$ (and similarly at $U/J=22$), $\bar{G}^{(2)}_{c}(\bm{0})/\Delta k^{3}$ follows a very different trend. While it may take slightly positive values at vanishingly small values of $\Delta k$ (where the amplitude is nonetheless compatible with being equal to zero), it is constant and negative otherwise. This different behaviour is at the origin of the negative values $\bar{G}^{(2)}_{c}(\bm{0})$ plotted in Fig. 2 of the main text. Observing $\bar{G}^{(2)}_{c}(\bm{0})<0$ implies that the number of coincidence counts in modes at opposite momenta is smaller than the coincidence counts expected in an uncorrelated ensemble of atoms with a similar momentum density $\langle N({\bm{k}})\rangle=\langle a^{\dagger}_{\bm{k}}a_{\bm{k}}\rangle$, and hence that modes at opposite momenta are anti-correlated.
We interpret the fact that $\bar{G}^{(2)}_{c}(\bm{0})/\Delta k^{3}$ is negative and constant for most values of $\Delta k$ (see Fig. S2 panel $U/J=20$) as an indication that the observed anti-correlation is not due to correlations between pairs of modes possessing a characteristic decay length. Indeed, two-mode correlations are generally expected to vanish beyond a given distance (in momentum space). A celebrated example of such a behavior is provided by the anti-bunching of ideal fermions Rom et al. (2006) which is present only at short distances. Instead when the anti-correlation originates from correlations between three bodies (or more), two-body correlation functions may not vary with the two-body distance.
III Normalised two-body correlations in the experiment
In the main text, we focus on connected correlations at opposite momenta, $G_{c}^{(2)}({\bm{k}},-{\bm{k}}+\delta{\bm{k}})=\langle N({\bm{k}})N({-\bm{k}}+\delta{\bm{k}})\rangle-\langle N({\bm{k}})\rangle\langle N({-\bm{k}}+\delta{\bm{k}})\rangle$. Alternatively, we could have used a normalised form of correlations defined as
$g^{(2)}(\bm{k},\bm{k}^{\prime})=\frac{\langle N({\bm{k}})N({\bm{k}}^{\prime})\rangle}{\langle N({\bm{k}})\rangle\langle N({\bm{k}}^{\prime})\rangle}.$ (S3)
In this section, we present our measurements of the normalised two-body correlation function $g^{(2)}(\bm{k},\bm{k}^{\prime})$ and we discuss the motivations to use non-normalised connected correlations $G_{c}^{(2)}(\delta{\bm{k}})$ to study the pairing strength.
In the weakly-interacting regime, at small values of $U/J$, the two-body correlation function $g^{(2)}(\bm{k},\bm{k}^{\prime})$ exhibits two peaks Tenart et al. (2021): one located at $\bm{k}^{\prime}=\bm{k}$ due to bosonic bunching (also referred to as Hanbury-Brown and Twiss (HBT) correlations) and one located at $\bm{k}^{\prime}=-\bm{k}$ due to interaction-induced quantum fluctuations. We are interested in extracting both the amplitudes of the bunching and of the pairing.
Proceeding as in Tenart et al. (2021) and along the lines described above in the case of $G_{c}^{(2)}({\bm{k}},{\bm{k}}^{\prime})$, we compute the terms $\langle N({\bm{k}})N({\bm{k}}^{\prime})\rangle$ and $\langle N({\bm{k}})\rangle\langle N({\bm{k}}^{\prime})\rangle$. We then obtain $g^{(2)}(\bm{k},\bm{k}^{\prime})$ from taking their ratio. From studying the variations of the normalised correlation functions with the volume used to compute them Tenart et al. (2021), we extrapolate the amplitudes, $g_{\rm HBT}^{(2)}(0)$ and $g_{\rm k/-k}^{(2)}(0)$, at zero transverse integration, with the results plotted in Fig. S3.
Figure S3: Magnitudes $g_{\rm HBT}^{(2)}(0)$ of local $k/k$ two-body correlations (bunching) and $g_{\rm k/-k}^{(2)}(0)$ of normalised two-body correlations at opposite momenta $k/-k$ plotted as a function of the interaction strength $u$ for data sets shown in Fig. 2 of the main text. Vertical error bars correspond to one standard deviation.
The bunching amplitude is perfectly-contrasted in all regimes of interaction, $g_{\rm HBT}^{(2)}(0)=2.00(5)$, extending previous observations Cayla et al. (2020) to the strongly-interacting regime. In contrast, $g_{k/-k}^{(2)}(0)$ is found to decrease with the interaction $u$. This decrease is compatible with the one expected in Bogoliubov theory: the magnitude $g_{k/-k}^{(2)}(0)-1$ is indeed expected to decrease with the inverse of the mode population $1/\langle N(\bm{k})\rangle$, a property of two-mode squeezed states Tenart et al. (2021).
Note that no qualitative change is expected when the gas enters the strongly correlated regime where the Bogoliubov approximation fails: $g_{k/-k}^{(2)}(0)$ also decreases when the population of two-mode squeezed states decreases. This illustrates that normalised correlation functions are not suited to study the hierarchy of momentum-correlated subsets introduced in the main text (see Fig. 1a) and it motivates the use of connected correlations to reveal the correlated subsets of $n$ modes.
IV Scaling of the uncertainty on the value of $G^{(2)}_{c}({\bf 0})$ with interactions
[…]
Incidentally, we notice that the error bars on the value of two-body connected correlations contain information about higher-order of correlations since $\left(\Delta\overline{G}_{c}^{(2)}(0)\right)^{2}=\sum_{k,k^{\prime}}\left[\langle N_{k}N_{-k}N_{k^{\prime}}N_{-k^{\prime}}\rangle-\langle N_{k}N_{-k}\rangle\langle N_{k^{\prime}}N_{-k^{\prime}}\rangle\right]$. In the future, it would be interesting to investigate whether the rapid increase reported in Fig. S4 reflects the contribution of these higher-order of correlations.
V Relation between $G^{(2)}_{c}({\bf 0})$ and four-point connected correlations $\langle a^{\dagger}_{\bf k}a^{\dagger}_{-\bf k}a_{\bf k}a_{-{\bf k}}\rangle_{c}$
We discuss the relation between the measured two-body connected correlations $G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ and four-point connected correlations $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}$ in the condensate depletion.
The four-point connected correlation (or cumulant) $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}$ is the difference between the four-point correlations $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle$ and all connected correlations of order less or equal to $n\leq 3$. For any four operators $X_{j|j=1..4}$, the four-point connected correlations can be expressed as a function of connected correlators of order $n\leq 3$,
$\langle X_{1}X_{2}X_{3}X_{4}\rangle_{c}$ $=$ $\langle X_{1}X_{2}X_{3}X_{4}\rangle$
-- $\Sigma_{c}^{3}$
-- $\Sigma_{c}^{2}$
-- $\langle X_{1}\rangle_{c}\langle X_{2}\rangle_{c}\langle X_{3}\rangle_{c}\langle X_{4}\rangle_{c}.$
where the term $\Sigma_{c}^{3}$ containing three-body connected correlations is
$\Sigma_{c}^{3}$ $=$ $\langle X_{1}X_{2}X_{3}\rangle_{c}\langle X_{4}\rangle_{c}+\langle X_{4}X_{1}X_{2}\rangle_{c}\langle X_{3}\rangle_{c}$
$+$ $\langle X_{3}X_{4}X_{1}\rangle_{c}\langle X_{2}\rangle_{c}+\langle X_{2}X_{3}X_{4}\rangle_{c}\langle X_{1}\rangle_{c}$
and that $\Sigma_{c}^{2}$ containing two-body connected correlations writes
$\Sigma_{c}^{2}$ $=$ $\langle X_{1}X_{2}\rangle_{c}\langle X_{3}X_{4}\rangle_{c}+\langle X_{1}X_{3}\rangle_{c}\langle X_{2}X_{4}\rangle_{c}$
$+$ $\langle X_{1}X_{4}\rangle_{c}\langle X_{2}X_{3}\rangle_{c}.$
Note that $\langle X_{j}\rangle_{c}=\langle X_{j}\rangle$.
In the condensate depletion, the average value of one operator is null, $\langle a_{\bm{k}}\rangle=\langle a^{\dagger}_{\bm{k}}\rangle=0$. Inserting this fact in the above-mentioned general expression of four-point connected correlations leads to,
[…]
$-|\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}\rangle|^{2}-|\langle a^{\dagger}_{\bm{k}}a_{-\bm{k}}\rangle|^{2}-\langle a^{\dagger}_{\bm{k}}a_{\bm{k}}\rangle\langle a^{\dagger}_{-\bm{k}}a_{-\bm{k}}\rangle.$
Since $G^{(2)}_{c}({\bm{k}},-{\bm{k}})=\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle-\langle a^{\dagger}_{-\bm{k}}a_{-\bm{k}}\rangle$ we obtain,